# Spectral Truncation Kernels: Noncommutativity in $C^{*}$-algebraic Kernel Machines

### v3

@@ 3.1 Convergence and interactions
We first show that $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$, $k^{\operatorname{prod},q}$, and $k^{\operatorname{sep},q}$ defined in Example 2 as $n$ goes to infinity, respectively. In this sense, the kernels $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are generalizations of typical kernels. In addition, these convergences imply that the interactions along $z\in\mathbb{T}$ in the kernels become smaller as $n$ grows since they disappear at $n=\infty$. In the following, all the proofs of theoretical statements are documented in Appendix A.
@@
For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$. If $x$ and $y$ are differentiable, then $k_{n}^{\operatorname{sep},q}(x,y)(z)\to k^{\operatorname{sep},q}(x,y)(z)$.
@@
Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 9 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq. (2). Lemma 10 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is by Brandolini & Travaglini (1997). We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq. (2), and apply Lemmas 9 and 10 to show Theorem 8. See Appendix A for more details.
As seen in Theorem 8, the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows. On the other hand, we also have the parameter $q$ that describes how many Toeplitz matrices created by $R_{n}$ are involved in the kernel. Regarding the relationship between $q$ and the interactions, we have the following remark.
Remark 11
We expect that the convergence in Theorem 8 becomes slower as $q$ becomes larger, and the interaction along $z\in\mathbb{T}$ becomes larger. This is based on the observation that the value related to the Fejér kernel on $\mathbb{T}^{q}$ is described by the corresponding value related to the Fejér kernel on $\mathbb{T}^{q-1}$, especially in the case of $q=2$. See Appendix B for more details.
@@ 3.2 Positive definiteness
To construct an RKHM, which is introduced in Subsection 2.1, the kernel $k$ has to be positive definite. Thus, we investigate the positive definiteness of the proposed kernels. Regarding $k_{n}^{\operatorname{poly},q}(x,y)$, we can show the positive definiteness since $x_{i}$ and $y_{i}$ are separated by the product. Regarding $k_{n}^{\operatorname{sep}}(x,y)$, since $x$ and $y$ depend only on $\tilde{k}$, we can use the positive definiteness of $\tilde{k}$.
@@
The kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{sep},q}$ are positive definite.
@@
We can set $\beta_{n}$ in $\hat{k}_{n}^{\operatorname{prod},q}$ as $n^{q}$ to guarantee the positive definiteness. However, even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite. In addition, considering non-positive kernels has also been investigated (Ong et al., 2004; Canu et al., 2005). Indeed, in practical computations in Subsection 7, the algorithms can work even if we set $\beta_{n}<n^{2q}$. Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
@@ 4 Generalization Bound
We now focus on what impact the truncation parameter $n$ has on generalization. We provide a generalization bound allowing us to observe the tradeoff between the representation power and the model complexity associated with the proposed kernels. According to Lemma 4.2 in Mohri et al. (2012), the generalization bound is described by the Rademacher complexity. We apply this lemma and derive a generalization bound.
Let $\Omega$ be a probability space with a probability measure $\mu$. Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be samples from a distributions of $\mathcal{A}_{0}^{d}$-valued random variable $X$ and $\mathcal{A}_{1}$-valued random variable $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$). Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$. Let $\mathrm{E}$ be the Bochner integral on $\Omega$ with respect to $\mu$. Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. We derive the following generalization bound for the kernels defined in Section 3.
@@
Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$, and $k_{n}^{\operatorname{sep},q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, and $D(k_{n}^{\operatorname{sep},q},x)=\tilde{k}(x,x)\prod_{j=1}^{q}\|R_{n}(a_{j})\|_{\operatorname{op}}^{2}$ for $x\in\mathcal{A}_{0}^{d}$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q},k_{n}^{\operatorname{sep},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
The first term of the right-hand side of (3) is the empirical loss, and the second term represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$. There is a tradeoff between these terms. If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ in $\mathcal{A}$ is larger, the outputs of ${k_{n}}$ can represent more functions, which enables us to describe the dependency of outputs on inputs more. Indeed, if $n=1$, then the output of $k_{n}$ is always a constant function, where the situation is the same as the existing separable kernel discussed in Section 1. Thus, the empirical loss (the first term) can become small if $n$ is large. On the other hand, according to Theorem 16, the complexity of $\mathcal{M}_{k_{n}}$ (the second term) is larger if $n$ is larger. An advantage of introducing the parameter $n$ to construct the new kernels based on the typical kernels in Example 2 is that we can control the empirical loss and the complexity through $n$. By setting a proper $n$, we can balance the two terms and make the expected loss $\mathrm{E}[g(f(X),Y)]$, the left-hand side of Eq. (3), small.
5 Application to Kernel Ridge Regression
We illustrate the effect of the proposed kernels by applying them to kernel ridge regression. Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ be input training samples and $y_{1},\ldots,y_{N}\in\mathcal{A}$ be output training samples. We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$, $k=k_{n}^{\operatorname{prod},q}$, or $k=k_{n}^{\operatorname{sep},q}$. Consider the typical minimization problem for regression, which is also considered by Hashimoto et al. (2023a):
@@ 5 Application to Kernel Ridge Regression
where $\mathbf{G}$ is the ${\mathcal{A}^{N\times N}}$-valued Gram matrix whose $(i,j)$-entry is defined as $k(x_{i},x_{j}){\in\mathcal{A}}$, $\mathbf{c}=[c_{1},\ldots,c_{N}]^{T}\in\mathcal{A}^{N}$, $\mathbf{y}=[y_{1},\ldots,y_{N}]^{T}\in\mathcal{A}^{N}$. Then, the solution $\mathbf{c}$ of the problem (5) satisfies $\mathbf{y}=(\mathbf{G}+\lambda I)\mathbf{c}$, which means $\mathbf{y}(z)=(\mathbf{G}(z)+\lambda I)\mathbf{c}(z)$. Therefore, for each $z\in\mathbb{T}$, we obtain $\mathbf{c}(z)$ by computing $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$.
@@
The computational cost for the product of two $n$ by $n$ Toeplitz matrices is $O(n^{2})$, and computing $S_{n}(A)(z)$ for an $n$ by $n$ matrix $A$ and $z\in\mathbb{T}$ costs $O(n^{2})$. Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$ or $\hat{k}_{n}^{\operatorname{prod},q}$ is $O((qn^{2}+mn^{2})N^{2})$. As for the computation of $\mathbf{c}(z)$, we need to compute $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$, which costs $O(N^{3})$. Thus, the total computational cost for obtaining $\mathbf{c}(z_{1}),\ldots,\mathbf{c}(z_{m})$ is $O((q+m)n^{2}N^{2}+mN^{3})$. Note that the cost is linear with respect to $m$. This is by virtue of considering function-valued kernels and RKHMs. The situation is different if we use vvRKHSs. For vvRKHSs, to obtain $m$ values as an output, we apply a $\mathbb{C}^{m\times m}$-valued kernel. If the kernel is nonseparable, the Gram matrix should be $mN$ by $mN$, which results in a computational cost of $O(m^{3}N^{3})$. We can reduce the cost with respect to $N$, e.g., by Nyström method (Drineas & Mahoney, 2005). Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel. However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.

### v4

@@ 3 $C^{*}$-algebra-valued Positive Definite Kernel with Spectral Truncation
3.1 Connection between the proposed kernel and existing kernels
The proposed function-valued kernels combined with the framework of RKHMs fill the gap between separable and commutative kernels with a computational cost lower than that of transformable kernels. The proposed kernels are separable if $n=1$ and commutative if $n=\infty$. Indeed, we have the following observation:
The case of $n=1$
The proposed kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ are equivalent to separable kernels:
$k_{1}^{\operatorname{poly},q}(x,y)(z)$ $=\sum_{i=1}^{d}\alpha_{i}\bigg{(}\int_{\mathbb{T}}\overline{x_{i}(t)}\mathrm{d}t\bigg{)}^{q}\bigg{(}\int_{\mathbb{T}}y_{i}(t)\mathrm{d}t\bigg{)}^{q},$
$k_{1}^{\operatorname{prod},q}(x,y)(z)$ $=\prod_{j=1}^{q}\int_{\mathbb{T}}\overline{\tilde{k}_{1,j}(x(t),y(t))}\mathrm{d}t\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),y(t))\mathrm{d}t.$
We can see that the output of the kernel does not depend on the variable $z$. For example, the output function with an input $x$ of the kernel ridge regression is in the form of $\sum_{i=1}^{N}k(x,x_{i})(z)c_{i}(z)$ for given samples $x_{1},\ldots,x_{N}\in\mathcal{X}$ and some $c_{1},\ldots,c_{N}\in C(\mathbb{T})$ (See Hashimoto et al. (2023a) or Section 5). Thus, the kernels identify dependencies between input and output functions separately. We can only capture global information of the input $x$ for obtaining the output.
The case of $n=\infty$
In the next section (Theorem 10), we will show that the proposed kernels $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ converge to commutative kernels $k^{\operatorname{poly},q}$ and $k^{\operatorname{prod},q}$ as $n$ goes to infinity. Thus, the proposed kernels are equivalent to commutative kernels:
$k_{\infty}^{\operatorname{poly},q}(x,y)(z)$ $=\sum_{i=1}^{d}\alpha_{i}(\overline{x_{i}(z)}y_{i}(z))^{q},$
$k_{\infty}^{\operatorname{prod},q}(x,y)(z)$ $=\prod_{j=1}^{q}\overline{\tilde{k}_{1,j}(x(z),y(z))}\tilde{k}_{2,j}(x(z),y(z)).$
The value of the output of the kernels at $z$ are determined only by $x(z)$ and $y(z)$. The kernels identify completely local dependencies between input and output functions.
The case of $1<n<\infty$
If $n$ is small, then we focus more on global information, and if $n$ is large, we focus more on local information. Indeed, as we discussed in Eq. (3), we have
[…]
where we assume $d=1$ for simplify the notation. The proposed kernels are described by the convolution of the input function and the Féjer kernel $F_{n}^{2q,P}$. If $n=1$, then $F_{n}^{2q,P}(z)=1$, and as $n$ goes to infinity, $F_{n}^{2q,P}$ goes to the delta function. This means that if $n$ is small, then the convolution with $F_{n}^{2q,P}$ extract global information more than local information. On the other hand, if $n$ is large, then the convolution with $F_{n}^{2q,P}$ extract local information more than global information. We illustrate how $F_{n}^{2,P}$ changes along $n$ in Figure 2.
Figure 2: Féjer kernel $F_{n}^{2,P}$ for $n=5,10,15$
Remark 8
We can determine optimal $n$ in the sense of the dependencies by observing the Féjer function $F_{n}^{2q,P}$. Since the proposed kernel is defined by the convolution of the input function with the Féjer kernel, the volume of the region where the value of the Féjer kernel is sufficiently large corresponds to the range of local dependencies. Thus, if we have a information of the local dependencies, then we can choose $n$ based on the values of $F_{n}^{2q,P}$. We also discuss a generalization of the proposed kernel to kernels with a continuous value parameter, instead of the discrete value parameter $n$, in Subsection 9.2. This generalization enables us to use an optimization method to learn a best possible parameter.
Remark 9
The proposed kernels are composed of complex-valued kernels $\tilde{k}_{i,j}$ in Definition 6. We can choose any kernel for $\tilde{k}_{i,j}$, and the properties of the proposed kernels depend also on that choice. If we choose a kernel with parameters as $\tilde{k}_{i,j}$ (such as a weighted sum of multiple kernels), then by optimizing the parameters, we can obtain a better kernel for given data or tasks. In addition, for function-valued kernels, in the same manner as the complex-valued kernels, the weighted sum of positive definite kernels and product of positive definite kernels are also a positive definite kernel. Thus, we can consider multiple proposed kernels and combine them with weight parameters. In that case, we can also optimize the parameters to obtain a better kernel for given data or tasks.
Computational cost
As we will discuss in the last paragraph of Section 5, the computational cost for the kernel regression task for obtaining an output function evaluated at $m$ points with the proposed kernel $k_{n}^{\operatorname{poly},q}$ or $k_{n}^{\operatorname{prod},q}$ is $O((q+m)n^{2}N^{2}+mN^{3})$. Thus, if $(q+m)n^{2}<m^{3}N$, then the proposed kernels are more computationally efficient than nonseparable and noncommutative kernels, such as the transformable kernels.
@@ 3.2 Convergence and interactions
We show theoretical results that support the discussion in Subsection 3.1. We first show that $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ defined in Definition 6 converge to $k^{\operatorname{poly},q}$ and $k^{\operatorname{prod},q}$ defined in Example 2 as $n$ goes to infinity, respectively. In the following, all the proofs of theoretical statements are documented in Appendix A.
@@
For $x,y\in\mathcal{A}^{d}$ and $z\in\mathbb{T}$, $k_{n}^{\operatorname{poly},q}(x,y)(z)\to k^{\operatorname{poly},q}(x,y)(z)$, $k_{n}^{\operatorname{prod},q}(x,y)(z)\to k^{\operatorname{prod},q}(x,y)(z)$ as $n\to\infty$.
@@
Since the product of Toeplitz matrices in $k_{n}^{\operatorname{poly},q}$ and $k_{n}^{\operatorname{prod},q}$ are represented by the sum over the indices in $\bigcup_{j=0}^{2q}Q_{j}^{m}$, Lemma 11 explains why the Fejér kernel with the polyhedron $P=\{r\in\mathbb{R}^{2q}\,\mid\,|\sum_{i=l}^{k}r_{i}|\leq 1,\ l\leq k\}$ appears in Eq. (3). Lemma 12 generalizes the convergence of the Fejér kernel on $\mathbb{T}$ to that on $\mathbb{T}^{q}$, which is proved by Brandolini and Travaglini (1997). We set $g(z)=\overline{x_{1}(z_{1})}\cdots\overline{x_{q}(z_{q})}y_{1}(z_{q+1})\cdots y_{q}(z_{2q})$ for $z=[z_{1},\ldots,z_{2q}]$, based on Eq. (3), and apply Lemmas 11 and 12 to show Theorem 10. See Appendix A for more details. Theorem 10 implies that the interactions along $z\in\mathbb{T}$ in the kernels become small as $n$ grows. This is because as $n$ goes to infinity, $F_{n}^{2q,P}$ goes to the delta function and by taking the convolution with the input function, it focuses more on local relationships between input and output functions as $n$ grows. Regarding the relationship between $q$ and the interactions see Appendix B.
@@ 3.3 Positive definiteness
To construct an RKHM, which is introduced in Subsection 2.1, the kernel $k$ has to be positive definite. Thus, we investigate the positive definiteness of the proposed kernels. Regarding $k_{n}^{\operatorname{poly},q}(x,y)$, we can show the positive definiteness since $x_{i}$ and $y_{i}$ are separated by the product.
@@
The kernel $k_{n}^{\operatorname{poly},q}$ is positive definite.
@@
We can set $\beta_{n}$ in $\hat{k}_{n}^{\operatorname{prod},q}$ as $n^{q}$ to guarantee the positive definiteness. However, even if we set a smaller $\beta_{n}$, $\hat{k}_{n}^{\operatorname{prod},q}$ may become positive definite. In addition, considering non-positive kernels has also been investigated (Ong et al., 2004; Canu et al., 2005). Indeed, in practical computations in Section 8, the algorithms can work even if we set $\beta_{n}<n^{2q}$. Deriving a tighter bound of $\beta_{n}$ or developing the theory for non-positive kernels for RKHMs is future work.
@@ 4 Generalization Bound
We now focus on what impact the truncation parameter $n$ has on generalization. According to Lemma 4.2 in Mohri et al. (2012), the generalization bound is described by the Rademacher complexity. We apply this lemma and derive a generalization bound.
Let $\Omega$ be a probability space with a probability measure $\mu$. Let $X_{1},\ldots,X_{N}$ and $Y_{1},\ldots,Y_{N}$ be samples from a distributions of $\mathcal{A}_{0}^{d}$-valued random variable $X$ and $\mathcal{A}_{1}$-valued random variable $Y$ on $\Omega$, respectively (i.e., for $z\in\mathbb{T}$, $X_{i}(z)$ is a sample from the distribution of $X(z)$). Here, $\mathcal{A}_{0}$ and $\mathcal{A}_{1}$ are subsets of $\mathcal{A}$. Let $\mathrm{E}$ be the Bochner integral on $\Omega$ with respect to $\mu$. Let $B>0$ and =$\mathcal{F}=\{f\in\mathcal{M}_{k}\,\mid\,\|f\|_{k}\leq B\}$. Let $g:\mathbb{R}\times\mathbb{R}\to\mathbb{R}_{+}$ be an error function. Assume there exists $L>0$ such that for $y\in\mathcal{A}_{1}$ and $z\in\mathbb{T}$, $x\mapsto g(x,y(z))$ is $L$-Lipschitz continuous. We derive the following generalization bound for the kernels defined in Section 3 based on Mohri et al. (2012); Hashimoto et al. (2023a) Here, to adapt to the generalization bound analysis, we assume the kernels are real-valued.
@@
Assume $k_{n}^{\operatorname{poly},q}$, $k_{n}^{\operatorname{prod},q}$ are real-valued. Let $D(k_{n}^{\operatorname{poly},q},x)=\sum_{j=1}^{d}\alpha_{j}\|R_{n}(x_{j})\|^{2q}_{\operatorname{op}}$, $D(\hat{k}_{n}^{\operatorname{prod},q},x)=\prod_{j=1}^{q}(\|R_{n}(\tilde{k}_{1,j}(x,x))\|_{\operatorname{op}}\|R_{n}(\tilde{k}_{2,j}(x,x))\|_{\operatorname{op}})+\beta_{n}C$, where $\|\cdot\|_{\operatorname{op}}$ is the operator norm and $C=\prod_{j=1}^{q}\int_{\mathbb{T}}\tilde{k}_{1,j}(x(t),x(t))\mathrm{d}t\int_{\mathbb{T}}\tilde{k}_{2,j}(x(t),x(t))\mathrm{d}t$. Assume $\beta_{n}\leq\beta_{n+1}$ for $\hat{k}_{n}^{\operatorname{prod},q}$. For $k_{n}=k_{n}^{\operatorname{poly},q},\hat{k}_{n}^{\operatorname{prod},q}$ and for any $\delta\in(0,1)$, with probability at least $1-\delta$, we have
@@
The first term of the right-hand side of (5) is the empirical loss, and the second term represents the model complexity of the RKHM $\mathcal{M}_{k_{n}}$. There is a tradeoff between these terms. If $n$ is larger, then since the approximation space $\operatorname{Span}\{e_{-(n-1)},\ldots,e_{n-1}\}$ in $\mathcal{A}$ is larger, the outputs of ${k_{n}}$ can represent more functions, which enables us to describe the dependency of outputs on inputs more. Thus, the empirical loss (the first term) can become small if $n$ is large. On the other hand, the complexity of $\mathcal{M}_{k_{n}}$ (the second term of the right-hand side of (5)) is larger if $n$ is larger.
5 Kernel Ridge Regression with the Spectral Truncation Kernels
We illustrate the effect of the proposed kernels by applying them to kernel ridge regression. Let $x_{1},\ldots,x_{N}\in\mathcal{A}^{d}$ be input training samples and $y_{1},\ldots,y_{N}\in\mathcal{A}$ be output training samples. We consider the case where we need the values of the output function evaluated at different $m$ points. We consider the RKHM $\mathcal{M}_{k}$ associated with $k=k_{n}^{\operatorname{poly},q}$ or $k=k_{n}^{\operatorname{prod},q}$. Consider the typical minimization problem for regression, which is also considered by Hashimoto et al. (2023a):
@@ 5 Kernel Ridge Regression with the Spectral Truncation Kernels
where $\mathbf{G}$ is the ${\mathcal{A}^{N\times N}}$-valued Gram matrix whose $(i,j)$-entry is defined as $k(x_{i},x_{j}){\in\mathcal{A}}$, $\mathbf{c}=[c_{1},\ldots,c_{N}]^{T}\in\mathcal{A}^{N}$, and $\mathbf{y}=[y_{1},\ldots,y_{N}]^{T}\in\mathcal{A}^{N}$. Then, the solution $\mathbf{c}$ of the problem (7) satisfies $\mathbf{y}=(\mathbf{G}+\lambda I)\mathbf{c}$, which means $\mathbf{y}(z)=(\mathbf{G}(z)+\lambda I)\mathbf{c}(z)$. Therefore, for each $z\in\mathbb{T}$, we obtain $\mathbf{c}(z)$ by computing $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$.
@@
The computational cost for the product of two $n$ by $n$ Toeplitz matrices is $O(n^{2})$, and computing $S_{n}(A)(z)$ for an $n$ by $n$ matrix $A$ and $z\in\mathbb{T}$ costs $O(n^{2})$. Thus, the cost for constructing $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$ for different $m$ points $z_{1},\ldots,z_{m}\in\mathbb{T}$ with $k_{n}^{\operatorname{poly},q}$ or $\hat{k}_{n}^{\operatorname{prod},q}$ is $O((qn^{2}+mn^{2})N^{2})$. As for the computation of $\mathbf{c}(z)$, we need to compute $(\mathbf{G}(z)+\lambda I)^{-1}\mathbf{y}(z)$, which costs $O(N^{3})$. Thus, the total computational cost for obtaining $\mathbf{c}(z_{1}),\ldots,\mathbf{c}(z_{m})$ is $O((q+m)n^{2}N^{2}+mN^{3})$. Note that the cost is linear with respect to $m$. This is by virtue of considering function-valued kernels and RKHMs. The situation is different if we use vvRKHSs. For vvRKHSs, to obtain $m$ values as an output, we typically apply a $\mathbb{C}^{m\times m}$-valued kernel. Then, the Gram matrix should be $mN$ by $mN$. Specifically, if the kernel is transformable, the computational cost is $O(m^{3}N^{3})$. Thus, if $(q+m)n^{2}<m^{3}N$, then the proposed kernels are more computationally efficient than the transformable kernels. For the proposed kernel, we can reduce the cost with respect to $N$, e.g., by applying Nyström method (Drineas and Mahoney, 2005) to each $\mathbf{G}(z_{1}),\ldots,\mathbf{G}(z_{m})$. Regarding $n$ and $q$, one approach to reducing the cost with respect to them is investigating the method to approximate the Fejér kernel. Generalization of the proposed kernel with the Fejér kernel to a kernel with another integral kernel, which we will discuss in Subsection 9.2 is also a way to reduce computational cost. However, the main contribution of this paper is to show the advantage of introducing the noncommutivity in kernels, and investigating computationally effective methods is future work.